What is the yield on a zero-coupon bond with a par value of $1,000, a price of $775 and a maturity of 6 years? Assume semi annual compounding.
Answer #1

As Semi Annual Compounding

Period = 6*2 = 12

YTM = (Maturity Value / Price)(1/n) - 1

   = (775 / 1000)(1/12) - 1

= 0.9790 - 1

YTM = 0.021 or 2.1%